OpenStudy (anonymous):

PLEASE HELP - GEOMETRIC SEQUENCE 2 QUESTIONS Graph the six terms of a finite series where a1 = -3 and r = 1.5 Graph the first six terms of a sequence where a1 = -10 and d = 3

OpenStudy (anonymous):

@girlover

OpenStudy (anonymous):

@TheAsker2002

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@oOKawaiiOo

OpenStudy (anonymous):

@just_one_last_goodbye

OpenStudy (anonymous):

@confusegirllll

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@Coolsector

OpenStudy (anonymous):

@AriPotta

OpenStudy (anonymous):

@HelpBlahBlahBlah

OpenStudy (anonymous):

@inkyvoyd

myininaya (myininaya):

geometric sequence looks like \[a_1,a_2=a_1r,a_3=a_1r^2,a_4=a_1r^3, ....\]

myininaya (myininaya):

arithmetic sequence looks like \[a_1a_2=,a_1+d,a_3=a_1+2d,a_4=a_1+3d,....\]